Ang Electronics, Inc., has developed a new DVDR. If the DVDR is successful, the present value of the payoff (when the product is brought to market) is $33.5 million. If the DVDR fails, the present value of the payoff is $11.5 million. If the product goes directly to market, there is a 50 percent chance of success. Alternatively, the company can delay the launch by one year and spend $1.25 million to test market the DVDR. Test marketing would allow the firm to improve the product and increase the probability of success to 80 percent. The appropriate discount rate is 11 percent. Calculate the NPV of going directly to market and the NPV of test marketing before going to market. (Do not round intermediate calculations and enter your answers in dollars, not millions of dollars, rounded to the nearest whole number, e.g., 1,234,567.)
Answer #1

Answer:

NPV of going directly to market = $22,500,000

NPV of test marketing before going to market = $24,966,216

Ang Electronics should do test marketing before going to market as this option has higher NPV

Working:

Going directly to market:

If successful, the present value of the payoff (when the product is brought to market) is = $33.5 million = $33,500,000

If the DVDR fails, the present value of the payoff is = $11.5 million = $11,500,000

There is a 50 percent chance of success.

Hence:

Probability of failure = 100% - 50% = 50%

NPV of going directly to market = 50% * 33500000 + 50% * 11500000 = $22,500,000

Test marketing before going to market:

Cost of test marketing = $1.25 million = $1,250,000

Test marketing would allow the firm to improve the product and increase the probability of success to 80 percent.

Probability of failure = 100% - 80% = 20%

Year 1 value = 80% * 33500000 + 20% * 11500000 = $29,100,000

NPV of test marketing before going to market = $29,100,000 /(1 + 11%) - 1250000 = $24,966,216.22